# Tuning-Free Alignment of Diffusion Models with Direct Noise Optimization

### v2

@@ 3.1 Quantifying Low-Probability Region via Concentration Inequalities
3.2 Probability-Regularized Noise Optimization (PRNO)
With the insights discussed above, a natural idea for preventing OOD reward hacking is to regularize noise vectors to remain within the high-probability region of the Gaussian distribution. To achieve this, we propose the following Probability-Regularized Noise Optimization (PRNO) problem:
@@ 4 Tackling Non-Differentiable Reward Functions
Method 3. We found that there is a crucial drawback in the gradient estimator (18), that one needs to query the reward model $r(\cdot)$ with noisy input $x+\mu\xi$. When the reward model is only defined on some manifold $\mathcal{M}$, e.g., defined on the image manifold, rather than the whole space ${\mathbb{R}}^{n}$, this can lead to severe problems, because, for some $x\in\mathcal{M}$, the noisy sample $x+\mu\xi$ may no longer stay within $\mathcal{M}$. To remedy this issue, we propose to perturb the sample through the latent noise, rather than directly in the sample space. Specifically, our proposed new gradient estimator for $\nabla r(x)$ is
@@
Following a similar proof in (Nesterov and Spokoiny, 2017), we can also show that $H_{2}(x)\approx C_{2}\nabla r(x)$ for some constants $C_{2}$. As we can see, when computing the gradient (20), we ensure that we query the reward model $r(\cdot)$ with only samples that are within the manifold of the pretrained distribution. Similar to Hybrid-1, we can adopt a gradient estimator $G_{2}(z)$ with $H_{2}(x)$:
@@ 5 Experiments
5.1 Examining the Effectiveness of Probability Regularization
Setting. In this section, we design an experiment to demonstrate the effectiveness of probability regularization proposed in Section 3.2. We consider two settings: The first involves optimizing the whiteness reward, i.e., the average value of all pixels—the larger this value is, the whiter the image becomes—with the prompt "black <animal>", where the token <animal> is randomly selected from a list of animals; The second setting involves optimizing the blackness reward, i.e., the negative of the whiteness reward, with the prompt "white <animal>".
The primary insight behind designing such experiments is the clear contradiction between the prompt and the reward, making it easier to trigger the OOD reward-hacking phenomenon. Moreover, it is straightforward to verify whether the generated samples is out-of-distribution by simply examining the color of the generated animals. In these experiments, we will compare the noise optimization process with and without probability regularization to understand the capability of probability regularization in preventing the OOD reward-hacking phenomenon.
Figure 7: Comparison of DNO with and without probability regularization. Upper row: Optimizing for the whiteness reward, i.e., the average value of all pixels in the images. Lower row: Optimizing for the blackness reward, i.e., the negative of the whiteness reward. The x-axis refers to the number of gradient ascent steps during optimization.
Measuring the Degree of OOD. As observed in this experiment, the reward function contradicts the input prompt, leading to inconsistency between the generated samples and the prompt. Therefore, we utilize the CLIP Score (CS) (Radford et al., 2021), a commonly used metric for measuring the semantic similarity between images and text descriptions, to gauge the degree of OOD for the generated samples. The higher the CS, the less likely the sample is to be out-of-distribution (OOD).
@@ 5.1 Examining the Effectiveness of Probability Regularization
Results. In Figure 7, we first observe that adding the regularization term leads to a mildly slower optimization process. However, the generated samples are much more consistent with the prompt throughout the entire optimization process, as reflected by the CS curve. The trajectory of $P(z)$ further corroborates that it is a good indicator of the OOD phenomenon, as it is positively associated with the CS. It is worth noting that the regularization effect provided by probability cannot be fully replicated by simply employing early-stopping in non-regularized optimization, as illustrated in Figure 7. We offer qualitative examples in Figure 8 to highlight the differences between optimized samples with and without regularization. Figure 8(a) and 8(b) are optimized images with whiteness reward and prompt "black rabbit". Figure 8(c) and 8(d) are optimized images with blackness reward and prompt "white chicken". In the two pairs of images, we select the optimized images that achieved the almost the same reward value from the optimizations with and without regularization, respectively. It is evident that at an equivalent level of reward, the non-regularized samples are more prone to be out-of-distribution, as the color of the generated animals is less likely to match the prompt.
5.2 Benchmarking Three Reward Models Trained on Human Feedback Data
Setting. In this section, we investigate the performance of the proposed method using three common reward models trained from human feedback data, specifically Aesthetic Score (Schuhmann et al., 2022b), HPS-v2 score (Wu et al., 2023), and PickScore (Kirstain et al., 2023), respectively. In this experiment, we also compare noise optimization with and without probability regularization. However, compared to the reward function used in the previous section, using these reward models presents a lesser chance for the optimized image to be OOD. In this case, to measure the benefit of probability regularization in maintaining the quality of the generated sample, we consider using the other two reward models as test metrics when optimizing one of them. For example, when optimizing the Aesthetic Score, we use the HPS Score and Pick Score as test metrics.
@@ 5.2 Benchmarking Three Reward Models Trained on Human Feedback Data
Figure 9: Comparison of running DNO with three human-like reward models, with and without regularization. When optimizing one reward model, the other two are used as test metrics. A, H, P are short for Aesthetic Score, HPS Score, and Pick Score, respectively. The name for each line comprises the used reward model and whether the regularization is used. For example, A + w/ reg means optimizing aesthetic score with regularization.
Comparing to Existing Alignment Methods. We also summarize the performance of our proposed method in Table 1 and compare it to the major existing alignment methods discussed in the introduction. As we can see, the proposed method can easily outperform state-of-the-art tuning-based alignment methods without any fine-tuning on the network models, within a reasonable time budget for generation. We found that another tuning-free method, LGD (Song et al., 2023a), performs badly with these complex reward functions, since it is impossible to estimate the gradient of reward functions without a complete generation process. Finally, we would like to highlight that this alignment process can run with memory usage of less than 15GB, and thus can easily fit into just one consumer-level GPU. In contrast, current tuning-based methods require significantly more computing resources, typically 4-8 advanced GPUs like A100 (Black et al., 2023; Prabhudesai et al., 2023).
Method SD v1.5 LGD SPIN DDPO AlignProp PRNO (This work)
@@
Aesthetic $\uparrow$ 5.367 5.340 5.224 6.248 7.180 8.940 5.754 7.202 8.587 10.68
HPS $\uparrow$ 0.278 0.276 0.271 0.276 0.287 0.330 0.285 0.303 0.324 0.354
PickScore $\uparrow$ 21.11 21.01 21.09 22.00 / / 21.25 23.17 25.13 25.16
Table 1: Performance comparison. For SD v1.5 and PRNO, we annotate the generation time below the name. For LGD, we annotate the number of samples used for Monte Carlo approximation. For SPIN, DDPO, and AlignProp, we annotate the estimated time for fine-tuning. All time costs in the table are measured with respect to the GPU time on a single A800 GPU. Baselines: LGD (Song et al., 2023a), SPIN (Yuan et al., 2024), DDPO (Black et al., 2023), AlignProp (Prabhudesai et al., 2023).
@@ 5.3 Optimization for Non-Differentiable Reward Functions
Setting. In this section, we aim to explore the three extensions proposed in Section 4 for handling non-differentiable reward functions. For Method 1, we use ZO-SGD (Nesterov and Spokoiny, 2017), and compare it with our proposed Hybrid-1 and Hybrid-2. We consider using two reward functions: The Jpeg Compressibility score, which is the file size of the image after compressing it using the JPEG algorithm. This reward model was also used in (Black et al., 2023) and is intrinsically non-differentiable. The second reward function is the Aesthetic Score used in the previous section. Unlike previous experiments, in this section we treat the Aesthetic Score as a non-differentiable reward function for optimization. The goal is to simulate a scenario where the neural network model of the Aesthetic Score can only be queried via an API provider that returns the score rather than its gradient. We compare the three methods in terms of optimization steps and the final score of the reward function. For ease of demonstration, we do not add the probability regularization term to the optimization process. Moreover, it is important to note that in these three methods, the major time expenditure comes from estimating the gradient with finite samples. For a fair comparison among these methods, we set the number of samples for estimating the gradient separately so that the total time spent estimating the gradient is the same for all three methods.
@@
Results. The main results are visualized in Figure 10. Initially, we observe that the Hybrid-2 method is significantly faster than the other two methods, and the final score is also higher for both reward functions used in this experiment. Furthermore, in both scenarios, the ZO-SGD method exhibits the slowest optimization speed. Another interesting finding is the poor performance of Hybrid-1 in optimizing the Aesthetic Reward. This mainly occurs because the aesthetic reward model can only work with image inputs, which validates our initiative to propose the Hybrid-2 method in Section 4. Finally, when comparing the optimization of the aesthetic score in Figures 9 and 10, we can also note that using the true gradient results in substantially fewer steps than using the estimated gradient.
@@ 6 Conclusions
In this work, we present a comprehensive study on Direct Noise Optimization (DNO) for aligning diffusion generative models. We introduce variants of DNO designed to efficiently address challenges such as out-of-distribution reward-hacking and the optimization of non-differentiable reward functions. More significantly, we demonstrate the exceptional efficacy of DNO, underscoring its capacity to rival tuning-based methods. The primary limitation of DNO lies in its integration with the sampling process of diffusion models, leading to a substantial increase in processing time compared to direct sampling. Nonetheless, we argue that the additional time cost, being within a reasonable and acceptable range, is a worthwhile trade-off for attaining high-reward generated samples in a wide range of real-world applications. We anticipate DNO gaining greater attention in future research and applications due to its flexibility to accommodate any reward function—or even a combination of different reward functions—while demanding only modest computing resources, which positions it as an accessible tool for many practitioners.
@@ A Visualization of Optimization with Images
(a) Optimizing whiteness reward with and without probability regularization. Prompt: "black duck".
(b) Optimizing blackness reward with and without probability regularization. Prompt: "white duck".
@@
Figure 12: Representative examples of optimizing reward models trained on human feedback data.
@@ A.2 More Examples
Firstly, in Figure 12, we present examples from the optimization of all three popular human-level reward models discussed in Section 5.2. As can be observed, the optimization process indeed results in an increase in human preference throughout.
@@ B Theoretical Results
Here, we denote $G_{1}(z_{0})=\nabla_{x}r(x)_{|x=M_{\theta}(g_{t}(z_{0}))}$ as the gradient of the reward models on the generated sample, and $G_{2}(z_{0})=\nabla_{z}M_{\theta}(z)_{|z=g_{t}(z_{0})}$ representing the Jacobian matrix of the noise-to-sample mapping. We categorize the situation in Equation (24) into three cases:
Type-I: $\|G_{1}(z_{0})\|=0$ and $\|G_{2}(z_{0})\|>0$. Here, the gradient of the reward model on the generated sample is zero, indicating that the generated sample has reached a stationary point (or local solution) of the reward model.
@@
Type-III: $\|G_{1}(z_{0})\|>0$ and $\|G_{2}(z_{0})\|>0$, but $\|G_{1}(z_{0})\cdot G_{2}(z_{0})|\ =0$. In this scenario, the gradient of the reward model on the generated sample is orthogonal to the Jacobian matrix of the noise-to-sample mapping.
@@
To showcase an example of Type-III noise vectors, we introduce a new toy example illustrated in Figure 14(b). Specifically, the ground-truth distribution learned by diffusion models is uniform across a horizontal line spanning from $(-1,0)$ to $(1,0)$. The reward function is defined as $r(x,y)=y$. It can be readily confirmed that, for every point on this line, the gradient of the reward model is orthogonal to the Jacobian matrix of the noise-to-sample mapping. Consequently, all points along the line segment $[(-1,0),(1,0)]$ qualify as Type-III noise vectors.

### v3

@@ 3.1 Quantifying Low-Probability Region via Concentration Inequalities
3.2 Probability-Regularized Noise Optimization
With the insights discussed above, a natural idea for preventing OOD reward hacking is to regularize noise vectors to remain within the high-probability region of the Gaussian distribution. To achieve this, we propose the following Probability-Regularized Noise Optimization problem:
@@ 4 Tackling Non-Differentiable Reward Functions
Method 3. We found that there is a crucial drawback in the gradient estimator (18), that one needs to query the reward function $r(\cdot)$ with noisy input $x+\mu\xi$. When the reward function is only defined on some manifold $\mathcal{M}$, e.g., defined on the image manifold, rather than the whole space ${\mathbb{R}}^{n}$, this can lead to severe problems, because, for some $x\in\mathcal{M}$, the noisy sample $x+\mu\xi$ may no longer stay within $\mathcal{M}$. To remedy this issue, we propose to perturb the sample through the latent noise, rather than directly in the sample space. Specifically, our proposed new gradient estimator for $\nabla r(x)$ is
@@
Following a similar proof in (Nesterov and Spokoiny, 2017), we can also show that $H_{2}(x)\approx C_{2}\nabla r(x)$ for some constants $C_{2}$. As we can see, when computing the gradient (20), we ensure that we query the reward function $r(\cdot)$ with only samples that are within the manifold of the pretrained distribution. Similar to Hybrid-1, we can adopt a gradient estimator $G_{2}(z)$ with $H_{2}(x)$:
@@ 5 Experiments
5.1 Experiments on Image Brightness and Darkness Reward Functions
Experiment Design. In this section, we design an experiment to demonstrate the effectiveness of DNO as described in Section 2, and the probability regularization proposed in Section 3.2. We consider two settings: The first involves optimizing the brightness reward, which is the average value of all pixels—the higher this value, the brighter the image becomes—with the prompt "black <animal>", where the token <animal> is randomly selected from a list of animals. The second setting involves optimizing the darkness reward, defined as the negative of the brightness reward, with the prompt "white <animal>". The primary rationale behind designing such experiments is the inherent contradiction between the prompt and the reward, which makes it easier to trigger the OOD reward-hacking phenomenon. Moreover, it is straightforward to verify whether the generated samples are out-of-distribution by simply examining the color of the generated animals. In these experiments, we compare the noise optimization process with and without probability regularization to assess the capability of probability regularization in preventing the OOD reward-hacking phenomenon.
Importance of the Brightness and Darkness Reward Functions. While the primary purpose of using these two reward functions is to better examine the effectiveness of probability regularization, it is also important to highlight their practical utility. There is often a genuine need to generate images with extremely dark or bright backgrounds, which cannot be achieved by the base models through prompting alone, as reported in the notable research by (CrossLabs, 2023).
Figure 6: Comparison of DNO with and without probability regularization. Upper row: Optimizing for the brightness reward, i.e., the average value of all pixels in the images. Lower row: Optimizing for the darkness reward, i.e., the negative of the brightness reward. The x-axis refers to the number of gradient ascent steps during optimization.
Measuring the Degree of OOD. As observed in this experiment, the reward function contradicts the input prompt, leading to an inconsistency between the generated samples and the prompt. Therefore, we utilize the CLIP Score (CS) (Radford et al., 2021), a commonly used metric for measuring the semantic similarity between images and text descriptions, to gauge the degree of OOD for the generated samples. A higher CS indicates that the sample is less likely to be out-of-distribution (OOD). In addition to the CS, we also use an MLLM-based score to measure the degree of OOD, specifically employing the Image-Text Matching (ITM) score from (Wang et al., 2024a) as the metric.
@@ 5.1 Experiments on Image Brightness and Darkness Reward Functions
Results. In Figure 6, we first observe that adding the regularization term leads to a mildly slower optimization process. However, the generated samples are much more consistent with the prompt throughout the entire optimization process, as reflected by the CS and ITM curves. The trajectory of $P(z)$ further corroborates that it is a good indicator of the OOD phenomenon, as it is positively associated with CS and ITM. We offer qualitative examples in Figure 7 to highlight the differences between optimized samples with and without regularization. Figure 7(a) and 7(b) are optimized images with brightness reward and prompt "black rabbit". Figure 7(c) and 7(d) are optimized images with darkness reward and prompt "white chicken". In the two pairs of images, we select the optimized images that achieved almost the same reward value from the optimizations with and without regularization, respectively. It is evident that at an equivalent level of reward, the non-regularized samples are more prone to be out-of-distribution, as the color of the generated animals is less likely to match the prompt.
Generating Images with Purely Bright/Dark Backgrounds. From the images demonstrated in Figure 7, we observe that DNO with regularization can lead to images with purely dark or bright backgrounds. It is worth noting that this is a remarkable result, as DNO requires no fine-tuning. Notably, as discussed in (CrossLabs, 2023), such an effect can only be achieved by fine-tuning the diffusion models using a technique called Offset-Noise.
5.2 Benchmarking on Three Human-Aligned Reward Functions
Setting. In this section, we investigate the performance of the proposed method using three common reward functions trained from human feedback data, specifically Aesthetic Score (Schuhmann et al., 2022b), HPS-v2 score (Wu et al., 2023), and PickScore (Kirstain et al., 2023), respectively. In this experiment, we also compare noise optimization with and without probability regularization. However, compared to the reward function used in the previous section, using these reward functions presents a lesser chance for the optimized image to be OOD. In this case, to measure the benefit of probability regularization in maintaining the quality of the generated sample, we consider using the other two reward functions as test metrics when optimizing one of them.
@@ 5.2 Benchmarking on Three Human-Aligned Reward Functions
Figure 8: Comparison of running DNO with three human-like reward functions, with and without regularization. When optimizing one reward function, the other two are used as test metrics. A, H, P are short for Aesthetic Score, HPS Score, and Pick Score, respectively. The name for each line comprises the used reward function and whether the regularization is used. For example, A + w/ reg means optimizing aesthetic score with regularization.
Comparison to Existing Alignment Methods. We summarize the performance of DNO with probability regularization from Figure 8 into Table 1 and compare it to the major existing alignment methods discussed in the introduction. As shown, the performance of DNO matches that of state-of-the-art tuning-based alignment methods without any fine-tuning on the network models, all within a reasonable time budget for generation. More importantly, we demonstrate that DNO provides an worthwhile trade-off between inference time and the reward of the generated samples. On one hand, another inference-time method, LGD (Song et al., 2023a), performs poorly with these complex reward functions, as it is impossible to estimate the gradient of the reward functions without a complete generation process. On the other hand, we also examined the most fundamental inference-time alignment algorithm, Best-of-N (BoN) Sampling, which generates N samples and selects the one with the highest reward. In this experiment, we fix the time budget for BoN to 10 minutes, and we observe that DNO outperforms it by a large margin, demonstrating that DNO presents a highly advantageous trade-off between inference time and reward.
Using Fewer Steps to Reduce Inference Time. In Figure 8, we show that our proposed DNO can achieve high reward values within approximately 3–5 minutes. While this already represents a highly advantageous performance, 3–5 minutes of optimization may still be prohibitively long in practice. However, we note that reducing the number of diffusion steps can significantly decrease optimization time. In Table 1, we fix the diffusion steps to 50, as this is the setting used by all other algorithms to ensure a fair comparison. In practice, 50 diffusion steps is a fairly long choice, and using 15–25 steps can still result in sufficiently good samples. Therefore, in Table 2, we present the performance of DNO with different numbers of denoising steps, fixing the time budget to 1 minute, as well as other hyperparameters like learning rate and reguralization coefficient. As Table 2 shows, with 15 diffusion steps, DNO can achieve high reward values within 1 minute.
Method SD v1.5 LGD SPIN DDPO AlignProp DNO (This work) BoN
@@
Aesthetic $\uparrow$ 5.367 5.340 5.224 6.248 7.180 8.940 5.754 7.202 8.587 6.531
HPS $\uparrow$ 0.278 0.276 0.271 0.276 0.287 0.330 0.285 0.303 0.324 0.298
PickScore $\uparrow$ 21.11 21.01 21.09 22.00 / / 21.25 23.17 25.13 22.09
Table 1: Performance comparison. For SD v1.5 and DNO, we annotate the generation time below the name. For LGD, we annotate the number of samples used for Monte Carlo approximation. For SPIN, DDPO, and AlignProp, we annotate the estimated time for fine-tuning. All time costs in the table are measured with respect to the GPU time on a single A800 GPU. Baselines: LGD (Song et al., 2023a), SPIN (Yuan et al., 2024), DDPO (Black et al., 2023), AlignProp (Prabhudesai et al., 2023).
$T$ 10 15 20 25 50
Aesthetic $\uparrow$ 6.992 7.496 6.773 6.381 5.754
HPS $\uparrow$ 0.342 0.341 0.306 0.293 0.285
PickScore $\uparrow$ 23.98 24.82 23.69 23.02 21.25
Table 2: Running DNO with different numbers of diffusion steps $T$, while fixing the time budget to 1 minute.
Finally, we would like to highlight that this alignment process can run with memory usage of less than 15GB, and thus can easily fit into just one consumer-level GPU. In contrast, current tuning-based methods require significantly more computing resources, typically 4-8 advanced GPUs like A100 (Black et al., 2023; Prabhudesai et al., 2023).
@@ 5.3 Optimization for Non-Differentiable Reward Functions
Setting. In this section, we aim to explore the three extensions proposed in Section 4 for handling non-differentiable reward functions. For Method 1, we use ZO-SGD (Nesterov and Spokoiny, 2017), and compare it with our proposed Hybrid-1 and Hybrid-2. We consider using two reward functions: The Jpeg Compressibility score, which is the file size of the image after compressing it using the JPEG algorithm. This reward function was also used in (Black et al., 2023) and is intrinsically non-differentiable. The second reward function is the Aesthetic Score used in the previous section. Unlike previous experiments, in this section we treat the Aesthetic Score as a non-differentiable reward function for optimization. The goal is to simulate a scenario where the neural network model of the Aesthetic Score can only be queried via an API provider that returns the score rather than its gradient. We compare the three methods in terms of optimization steps and the final score of the reward function. For ease of demonstration, we do not add the probability regularization term to the optimization process. Moreover, it is important to note that in these three methods, the major time expenditure comes from estimating the gradient with finite samples. For a fair comparison among these methods, we set the number of samples for estimating the gradient separately so that the total time spent estimating the gradient is the same for all three methods.
@@
Results. The main results are visualized in Figure 9. Initially, we observe that the Hybrid-2 method is significantly faster than the other two methods, and the final score is also higher for both reward functions used in this experiment. Furthermore, in both scenarios, the ZO-SGD method exhibits the slowest optimization speed. Another interesting finding is the poor performance of Hybrid-1 in optimizing the Aesthetic Reward. This mainly occurs because the aesthetic reward function can only work with image inputs, which validates our initiative to propose the Hybrid-2 method in Section 4. Finally, when comparing the optimization of the aesthetic score in Figures 8 and 9, we can also note that using the true gradient results in substantially fewer steps than using the estimated gradient.
@@ 6 Conclusions
In this work, we present a comprehensive study on Direct Noise Optimization (DNO) for aligning diffusion generative models at inference-time. We introduce variants of DNO designed to efficiently address challenges such as out-of-distribution reward-hacking and the optimization of non-differentiable reward functions. More significantly, we demonstrate the exceptional efficacy of DNO, underscoring its capacity to rival tuning-based methods. The primary limitation of DNO lies in its integration with the sampling process of diffusion models, leading to a substantial increase in processing time compared to direct sampling. Nonetheless, we argue that the additional time cost, being within a reasonable and acceptable range, is a worthwhile trade-off for attaining high-reward generated samples in a wide range of real-world applications. We anticipate DNO gaining greater attention in future research and applications due to its flexibility to accommodate any reward function—or even a combination of different reward functions—while demanding only modest computing resources, which positions it as an accessible tool for many practitioners.
@@ A Visualization of Optimization with Images
(a) Optimizing brightness reward with and without probability regularization. Prompt: "black duck".
(b) Optimizing darkness reward with and without probability regularization. Prompt: "white duck".
@@
Figure 11: Representative examples of optimizing reward functions trained on human feedback data.
@@ A.2 More Examples
Firstly, in Figure 11, we present examples from the optimization of all three popular human-level reward functions discussed in Section 5.2. As can be observed, the optimization process indeed results in an increase in human preference throughout.
@@ B Theoretical Results
Here, we denote $G_{1}(z_{0})=\nabla_{x}r(x)_{|x=M_{\theta}(g_{t}(z_{0}))}$ as the gradient of the reward functions on the generated sample, and $G_{2}(z_{0})=\nabla_{z}M_{\theta}(z)_{|z=g_{t}(z_{0})}$ representing the Jacobian matrix of the noise-to-sample mapping. We categorize the situation in Equation (24) into three cases:
Type-I: $\|G_{1}(z_{0})\|=0$ and $\|G_{2}(z_{0})\|>0$. Here, the gradient of the reward function on the generated sample is zero, indicating that the generated sample has reached a stationary point (or local solution) of the reward function.
@@
Type-III: $\|G_{1}(z_{0})\|>0$ and $\|G_{2}(z_{0})\|>0$, but $\|G_{1}(z_{0})\cdot G_{2}(z_{0})|\ =0$. In this scenario, the gradient of the reward function on the generated sample is orthogonal to the Jacobian matrix of the noise-to-sample mapping.
@@
To showcase an example of Type-III noise vectors, we introduce a new toy example illustrated in Figure 13(b). Specifically, the ground-truth distribution learned by diffusion models is uniform across a horizontal line spanning from $(-1,0)$ to $(1,0)$. The reward function is defined as $r(x,y)=y$. It can be readily confirmed that, for every point on this line, the gradient of the reward function is orthogonal to the Jacobian matrix of the noise-to-sample mapping. Consequently, all points along the line segment $[(-1,0),(1,0)]$ qualify as Type-III noise vectors.